The most recent financial statements for Bello Co. are shown here: Income Statement Balance Sheet Sales $ 19,700 Current assets $ 11,860 Debt $ 16,180 Costs 13,440 Fixed assets 29,700 Equity 25,380 Taxable income $ 6,260 Total $ 41,560 Total $ 41,560 Taxes (25%) 1,565 Net income $ 4,695 Assets and costs are proportional to sales. Debt and equity are not. The company maintains a constant 30 percent dividend payout ratio. What is the sustainable growth rate? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

ROE=net income/equity

=4,695/25,380

=0.184988179

Retention ratio=1-payout ratio

=(1-0.3)=0.7

Sustainable growth rate=(ROE*Retention ratio)/[1-(ROE*Retention ratio)]

=(0.184988179*0.7)/[1-(0.184988179*0.7)]

=14.88%(Approx).